OpenStudy (gabylovesyou):

@waterineyes Determine if triangle ABC with coordinates A (−1, 1), B (1, −1), and C (3, 3) is an isosceles triangle. Use evidence to support your claim. If it is not an isosceles triangle, what changes can be made to make it isosceles? Be specific. I just did this and i got that its NOT isosceles.. correct ?

OpenStudy (gabylovesyou):

i got sqrt(8) sqrt(20) sqrt(-20)

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

i'm picasso btw

OpenStudy (anonymous):

Again find the length of AC.. :)

OpenStudy (anonymous):

you should not get -ve number in square root, this is just not possible..

OpenStudy (anonymous):

A(-1,1) and C(3,3) \[AC = \sqrt{(3-1)^2 + (3-(-1))^2} = \sqrt{(2)^2 + (3 +1)^2} = \sqrt{4 + 16} = ??\]

OpenStudy (anonymous):

Not getting @Gabylovesyou ??

OpenStudy (gabylovesyou):

but A is x2 and y2

OpenStudy (anonymous):

yeah, solve it further.

OpenStudy (anonymous):

your forgot to take sqrt of whole thing.. :)

OpenStudy (anonymous):

side CA sqrt{(-1 - 3)^2 + (1 - 3)^2 }

OpenStudy (anonymous):

then how you got sqrt{-20} there? :P

OpenStudy (anonymous):

So, two sides are equal or not?

OpenStudy (gabylovesyou):

idk.. wait but first is sqrt8 right /

OpenStudy (anonymous):

second is?

OpenStudy (gabylovesyou):

sqrt 20

OpenStudy (anonymous):

and third is

OpenStudy (gabylovesyou):

sqrt 20

OpenStudy (anonymous):

so two sides are equal no?

OpenStudy (gabylovesyou):

yep .. ok thanks

OpenStudy (anonymous):

any two sides should be equal, if third side is also equal then it is not isosceles but equilateral triangle in that case, getting?

OpenStudy (gabylovesyou):

yes

OpenStudy (anonymous):

That is great..!! :)